# Displacement Memory Effect from Supersymmetry

### v1

@@ 6 Discussion
In this work, based on the idea that DME conditions in plane GWs match the ones in quantum mechanics, we rederive the recent results of [9, 10, 11] within the supersymmetric framework. However, our work extends a mere rederivation in several ways. First of all, SUSY methods unify the Scarf and Pöschl-Teller GW profiles in an elegant way, showing that they are two different reductions of the same theory. The reduction process amounts to satisfying a couple of equations, which automatically restores the amplitudes of the wave profiles: no fine-tuning is necessary.
Geodesics with DME, i.e., $\Delta X\neq 0$, correspond to certain bound states of analogous quantum problems. Even in that case, the explanation depends on the parity and the quantization of the parameter in a certain way. Thanks to the shape invariance property of SUSY QM, solutions for the geodesic motion can be found readily using Darboaux transformations. For the PT profile, DME occurs with odd $n$. For the Scarf profile, DME occurs for all $n$. See Table 1 for some of our results.
@@
In addition, we propose a hypothetical 2-d transverse PT plane wave, which explains its half DME in itself. $4$-dimensional Scarf exact GW fits to our framework because of the $\pm$ ambiguity in its quantized parameter $B$.
Lastly, we make a first attempt to study DME in GWs with $\frac{1}{U^{2}}$ singularity. Our SUSY motivated model predicts the existence of the DME, see Section 5. Because of singularity, this investigation requires much care, and we reserve its complete treatment for a follow-up study where we also plan to work out DME in cosmological Lukash GW.
Another research problem is to look for DME in the Kepler GW [32], which corresponds to the Coulomb potential in quantum mechanics. Thanks to the integrability, we know the solutions of the Sturm-Liouville problem which may allow us to go to BJR coordinates and write down the Carroll symmetry generators at once.
There are further questions to be explored. The list of SUSY QM provides us with other exactly solvable potentials. Among them, there are trigonometric Scarf and trigonometric Pöschl-Teller potentials which can be prototypes for periodic GWs. We would like to study DME of periodic GWs using them. Another interesting direction is the double copy of the Scarf and the PT profiles that we worked out. We wonder what kind of singular electromagnetic configurations they yield and what their helicities will be.
As we know, the Pöschl-Teller equation frequently occurs in soliton physics. We would like to understand whether there is a relation between the DME of PT profile and the topology of solitons. Likewise, it would be interesting to understand the role of the Schwarzian derivative in [12] in the Scarf case. Finally, there is a recent memory type called the frequency memory effect [21] in plane GWs. We can also study such effects in our SUSY framework in the future.
@@ Acknowledgments
We are grateful to Krzysztof Andrzejewski, Peter Horvathy and Zurab Silagadze for their comments on our work.
@@
Appendix A SUSY QM: A brief review
Below, we provide a brief summary of supersymmetric methods in QM to be self-consistent. We suggest [5, 6] for more details.
We begin with the main elements of SUSY QM, namely its superpotential $W(U)$ and ladder operators $L,L^{\dagger}$ such that
$L=\frac{d}{dU}+W(U),\quad L^{\dagger}=-\frac{d}{dU}+W(U).$ (A.1)
They factorize Hamiltonians
$H_{1}$ $=L^{\dagger}L=-\frac{d^{2}}{dU^{2}}+V_{1},\quad\quad V_{1}=W^{2}-\frac{dW}{dU},$ (A.2a)
$H_{2}$ $=LL^{\dagger}=-\frac{d^{2}}{dU^{2}}+V_{2},\quad\quad V_{2}=W^{2}+\frac{dW}{dU},$ (A.2b)
where $V_{1,2}$ are partner potentials sharing almost the same spectrum, except the zero-energy ground state of $V_{1}$. SUSY also relates the scattering states of that partner potentials. If the potential is shape invariant, the exact spectrum can be found easily. It is this crucial property of shape invariance of SUSY QM that will allow us to derive DME of plane GWs.
The time-independent Schrödinger equation for the first potential $V_{1}$ is
$\left(-\frac{d^{2}}{dU^{2}}+V_{1}\right)\psi_{n}=E_{n}\psi_{n},$ (A.3)
where $n=0,1,2,...$ indexes the bound state. Up to a normalization factor, zero energy bound state $\psi_{0}(U)$ can be found via
[…]
immediately.
If partner potentials $V_{1,2}$ are shape invariant, i.e. connected via
$V_{2}(U;A_{1})=V_{1}(U;A_{2})+f(A_{1}),$ (A.5)
where $A_{1,2}$ are parameters [^7] Recall that plane wave profiles (2.4), (2.6) and (2.7) also have parameters like $k,g,..$., energy eigenvalues and eigenstates of (A.2) can be computed easily. Similar to $f(A_{1})$, $A_{2}$ is a function of the first parameter as well, i.e., $A_{2}=A_{2}(A_{1})$. Due to (A.5), discrete eigenstates are found step by step
$\psi_{n}(U;A_{1})=L^{\dagger}(U;A_{1})\psi_{n-1}(U;A_{2}).$ (A.6)
For instance, if $A_{1}$ and $A_{2}$ are related by a simple relation $A_{1}-A_{2}=1$, the excited state wave functions can be computed with subsequent action of $L^{\dagger}$ (A.1)
$\psi_{n}(U;A)=L^{\dagger}(U;A)L^{\dagger}(U;A-1)...L^{\dagger}(U;A-n+1)\psi_{0}(U;A-n),$ (A.7)
where we set $A_{1}=A$. In that case, (A.6) simplifies as
$\psi_{n}(U;A)=\left(-\frac{d}{dU}+W(U;A)\right)\psi_{n-1}(U;A-1).$ (A.8)
For now, the parameter $A$ can take any value. It has a physical meaning when applied to geodesics at Sections 3, 4, and 5.

### v3

@@ 6 Discussion
In this work, we reveal a hidden supersymmetry in the DME problem of particular plane waves and recover the recent results of [14, 15, 16] within the supersymmetric framework. However, our study goes beyond a mere rederivation in several ways. First of all, SUSY methods unify the Scarf and Pöschl-Teller GW profiles in an elegant way, showing that they are two different reductions of a common ancestor (3.9). The reduction process amounts to satisfying a couple of equations either (3.19) or (3.31), which automatically yields the amplitudes of the wave profiles: no fine-tuning is necessary. For PT profile, this reduction quantizes the shape invariance parameter $A$ and sets $B=0$. Therefore, we are left with a family of profiles labeled by $n\geq 1$. Due to the parity arguments, geodesics only with odd $n$ have a non-vanishing DME. When it comes to the Scarf case, we observe that the parameter $B$ is quantized non-trivially with a sign ambiguity (3.32). As a result, one may build a $4$-dimensional exact plane GW (4.1) using this property which is absent in the PT case. Thanks to shape invariance, the complicated solutions for the geodesics can be found readily using Darboux transformations. See Table 1 for some of our results.
@@
In addition, we propose a hypothetical 2-d transverse PT plane wave, which explains its half DME. $4$-dimensional Scarf exact GW fits to our framework because of the sign ambiguity in its quantized parameter $B$ (3.32).
Lastly, we study DME in GWs with $U^{-2}$ singularity. Our SUSY motivated model (5.3) predicts the existence of the DME for such type of GWs. Because of the singularity, this investigation requires much care, and we reserve its complete treatment for a follow-up study where we also plan to work out DME in cosmological Lukash GW.
Another research problem is to look for DME in the GW originating from the Kepler potential [38]. Thanks to SUSY, we may determine the solutions of the related SL problem that leads us to BJR coordinates. Based on this, we can compare the Carroll symmetry generators of plane GWs with the symmetries of the associated quantum mechanical problem.
There remain further questions to be explored. The list of SUSY QM provides us with other exactly solvable potentials. Among them, there are trigonometric Scarf and trigonometric Pöschl-Teller potentials which can be prototypes for periodic GWs [18]. We would like to study DME of periodic GWs using them. Another interesting direction is the double copy [39] of the Scarf (4.1) and of the PT (4.5) profiles that we have worked out. We wonder what kind of singular electromagnetic configurations they yield and what their helicities will be.
So far we studied the geodesic equations of spinless test particles (2.2) based on that the DME conditions of plane waves match the ones in non-relativistic Schroedinger theory. However, there is also supersymmetry in Pauli and Dirac equations [40, 41, 42]. Thus, our studies might be extended to trajectories of spinning particles along those arguments.
As is known, Pöschl-Teller equation frequently occurs in soliton physics. We would like to understand whether there is a relation between the DME of PT profile and the topology of solitons. Likewise, it would be interesting to understand the role of the Schwarzian derivative in [17] for the Scarf GW. Finally, there is a recent type of memory called the frequency memory effect [27] in plane GWs. We intend to study such effects in our SUSY framework in the future.
@@ Acknowledgments
We are grateful to Krzysztof Andrzejewski, Janos Balog, Peter Horvathy, Cem Yetişmişoğlu and Zurab Silagadze for their comments.
@@
Data Availability Statement: No data associated in the manuscript.